OpenStudy (anonymous):

2x-(x+10)=45 need help with this problem

OpenStudy (anonymous):

What do you need to do with the equation? Solve for x?

OpenStudy (anonymous):

removing the brackets gives 2x-x-10=45 as you were subtracting everything inside the bracket. this simplifies to x-10=45, and then adding 10 to both sides gives x=55

OpenStudy (anonymous):

Another way of doing this is to add (x+10) to both sides to give 2x = 45 + (x+10) 2x = x +55 Then subtract x from both sides. x = 55

OpenStudy (anonymous):

answer to: 15x=3x+120 and: 3x+(x-8)(=120)

OpenStudy (anonymous):

Combine all the x's into a single term. Do you know the commutative, distributive and associative laws?